OpenStudy (anonymous):

The sum of two numbers is 30. One number is twice as large as the second number.

OpenStudy (anonymous):

One number can be represented as x and the other number can be represented as 2x since it's twice as large as the other. In an equation it's 2x + x = 30 which is 3x = 30 then solve for x and you get your value as 10

OpenStudy (gorv):

let a+b=30 a=2b 2b+b=30 b=10 a=20